OpenStudy (anonymous):

Part 1: Find the tangent line approximation to cos x at x=π/4. Part 2: To one decimal place, estimate the maximum value of the error on the interval 0≤x≤π/2. please explain! Thanks :)

OpenStudy (whpalmer4):

Do you know how to find the tangent of \(\cos x\) at a given point?

OpenStudy (anonymous):

umm not sure :/ do you mean the tangent equation at cos x when x=pi/4? if so, would it be what it says in this link? http://www.wolframalpha.com/input/?i=tangent+equation+of+cos+x+at+x%3Dpi%2F4

OpenStudy (whpalmer4):

Yes, that's what I mean. As I recall, you're taking a calculus class, right?

OpenStudy (anonymous):

yes i am :) and for this, it's linear approximation, so i'd be using the formula f(a+∆x) roughly equals f(a)+f'(a)(x-a)?

OpenStudy (whpalmer4):

Yes.

OpenStudy (anonymous):

so how would I find part 1 and part 2? via linear approximation?

OpenStudy (whpalmer4):

You start by finding \(f'(x)\)

OpenStudy (anonymous):

of the equation cos x? or of the result from that link?

OpenStudy (whpalmer4):

\[f(x) = \cos x\]\[f'(x) =\]

OpenStudy (anonymous):

-sin x?

OpenStudy (whpalmer4):

You are trying to approximate \(f(x) = \cos x\). To do that, you need to find the derivative, which gives you the slope of the tangent line at any point.

OpenStudy (whpalmer4):

Yes. Now, what does \(-\sin(x) =\) when \(x = \pi/4\)?

OpenStudy (anonymous):

ahh okay... so -sin(pi/4) = -1/(√2) ?

OpenStudy (whpalmer4):

Yes, it does. So that's the slope of the tangent line to \(\cos(x)\) at \(x=\pi/4\). Now, we can do things from first principles, and construct the equation of the line through that point with that slope, and use that as our approximation, or we can use the formula which has already done so.

OpenStudy (anonymous):

oh yay! so do we use y-y1=m(x-x1) now? :/ or are we using f(a+∆x) = f(a)+f'(a)(x-a)?

OpenStudy (whpalmer4):

Well, those are the two choices. However, if you think about it, they are equivalent.

OpenStudy (whpalmer4):

More a matter of how you understand the situation — you get the same answer both ways.

OpenStudy (anonymous):

ohh okay.. in this case, which would you recommend?

OpenStudy (whpalmer4):

I tend to remember the basic constructs and re-derive more complicated ones as necessary, rather than trust my memory to be 100% accurate. So in this case, I would probably go the former route, and then say "oh yeah" after doing so and realizing that I could have remembered the other. But really, they amount to the same thing.

OpenStudy (anonymous):

okay so we have y-0=(-1/√2)(x-(pi/4)) ?

OpenStudy (whpalmer4):

So, we know that \(x_1 = \pi/4\) and \(y_1 = \cos(x_1) = \cos(\pi/4) = \frac{1}{\sqrt{2}}\) \[y - \frac{1}{\sqrt{2}} = -\frac{1}{\sqrt{2}}(x-\frac{\pi}{4})\]\[y = -\frac{1}{\sqrt{2}}x + \frac{\pi}{4\sqrt{2}} + \frac{1}{\sqrt{2}}\] which is exactly what WolframAlpha gave us.

OpenStudy (anonymous):

ohhh okay so that is the tangent line approximation? part 1 is: what you wrote above?

OpenStudy (whpalmer4):

so our approximation could be written as \[y = g(x) = -\frac{1}{\sqrt{2}}x + \frac{\pi}{4\sqrt{2}} + \frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

ahhh okay haha answered my question :P woo! so part 2 now?

OpenStudy (whpalmer4):

Here's a graph, I trust you can figure out which line is which!

OpenStudy (whpalmer4):

Now we want to estimate the maximum error over that region. Any idea how we can compute the error at any given point?

OpenStudy (anonymous):

the red one right? ermm not sure... plug a number in?

OpenStudy (whpalmer4):

what is the error? What does that term mean to you?

OpenStudy (anonymous):

would that mean what is the exact equation/number/solution? because what we just got was the approximation, so would we have to find the exact and see what the difference is or something like that?

OpenStudy (whpalmer4):

Yes, the function - the approximation gives you the error.

OpenStudy (anonymous):

okay :) how can i find that ? do i use that equation we got above and plug in pi/4 wherever there's an x?

OpenStudy (whpalmer4):

You're not getting this, are you? \[Error(x) = f(x) - g(x)\]where \(g(x)\) is the function that we determined for our approximation. At \(x=\pi/4\), the error is 0. Everywhere else on that graph, the error is some number other than 0.

OpenStudy (whpalmer4):

Here's a graph that includes the error.

OpenStudy (anonymous):

yeah calc kinda confuses me :( but i see now.. so g(x) was what we found earlier and f(x) is the exact equation?

OpenStudy (anonymous):

oh the error is the curve at the bottom?

OpenStudy (whpalmer4):

Yeah.

OpenStudy (anonymous):

okay:) so how would we find the max value rounded to one decimal place? because it has to be an exact number right?

OpenStudy (whpalmer4):

First, you need to figure out where the error is going to be a maximum, yes?

OpenStudy (whpalmer4):

When you know where, then you can calculate the actual error.

OpenStudy (anonymous):

is it from 0.5-1?

OpenStudy (whpalmer4):

Even if you ignore the graph where I showed the error, it shouldn't be too terribly difficult to figure out the value of \(x\) where our approximation is furthest from the actual value of \(\cos x\)

OpenStudy (anonymous):

pi/4√2 ?

OpenStudy (whpalmer4):

what's that? your shoe size?

OpenStudy (whpalmer4):

Look at the graph!

OpenStudy (anonymous):

hahaha :P at -0.2 ?

OpenStudy (whpalmer4):

Where are the purple and blue lines close together? Where are they far apart?

OpenStudy (anonymous):

1.0 and 1.2 ?

OpenStudy (anonymous):

^farthest apart

OpenStudy (whpalmer4):

My graph starts at \(x = 0\). It goes to \(x = \pi/2 \approx 1.5708\). At what value of \(x\) are the blue and purple lines separated by the largest horizontal distance?

OpenStudy (whpalmer4):

That separation is the value of the error at that point.

OpenStudy (whpalmer4):

sorry, vertical distance, not horizontal distance.

OpenStudy (anonymous):

atx=0 ? not sure if i'm saying that right... but this is where i see the largest vertical distance... is that right?

OpenStudy (whpalmer4):

yes. at x = 0. Now, estimate, to 1 decimal place, the error. The convention is that you take the actual value of the function and subtract the approximated value to get the error.

OpenStudy (anonymous):

0.2 ?

OpenStudy (whpalmer4):

What is the value of \(\cos 0 =\)

OpenStudy (anonymous):

cos0=1 ?

OpenStudy (whpalmer4):

You tell me.

OpenStudy (whpalmer4):

I'll wait until you're sure.

OpenStudy (anonymous):

okay yeah cos 0 = 1 :)

OpenStudy (whpalmer4):

I may wander off and go to bed, but I'll wait :-)

OpenStudy (anonymous):

haha :P

OpenStudy (whpalmer4):

Okay. What is the value of the approximation at x = 0?

OpenStudy (anonymous):

thanks!! yeah sorry this is taking so long!! i'm not understanding this too well :/ more practice and studying is needed!! i got 1.817827516.... is that right?

OpenStudy (anonymous):

so the error is 0.8 ?

OpenStudy (whpalmer4):

No. Show me how you got that number. Does it look like the value from the graph?

OpenStudy (whpalmer4):

\[y = g(x) = \frac{1}{\sqrt{2}} + \frac{\pi}{4\sqrt{2}} - \frac{x}{\sqrt{2}}\]\[g(0) = \frac{1}{\sqrt{2}} + \frac{\pi}{4\sqrt{2}} - \frac{0}{\sqrt{2}} \approx \frac{1}{1.414} + \frac{3.14}{4*1.414} =\]

OpenStudy (anonymous):

no :/ umm \[-\frac{ 1 }{ \sqrt{2} }(0)+\frac{ \pi }{ 4\sqrt{2} }+\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (anonymous):

ohh i accidentally plugged the wrong way.. oops so from there you should actually get 0.5551626591 ? so the error is (when rounded to one decimal place) 0.6?

OpenStudy (whpalmer4):

So that's about 0.707 + 3.14/5.656

OpenStudy (anonymous):

oh so you get 1.262162659 and 1.262162659-1 = 0.2621626591 so the error is about 0.3 ?

OpenStudy (whpalmer4):

This is an example of how the proliferation of calculators has destroyed the good number sense of today's students...and then if you can't operate the calculator accurately, you not only don't get the right answer, but you don't even have a sense of whether you should be questioning whether you got the right answer! :-(

OpenStudy (whpalmer4):

Once again. Take the actual value of the function. Subtract the approximated value. What do you get?

OpenStudy (anonymous):

yeah that and poor understanding of the subjects lol i got 0.26216 :/ is that right?

OpenStudy (whpalmer4):

What is the actual value of the function?

OpenStudy (whpalmer4):

What is the approximated value of the function?

OpenStudy (whpalmer4):

What is the actual value of the function - the approximated value of the function?

OpenStudy (anonymous):

actual value is 1.26247 approx value is 1.26216 ?

OpenStudy (whpalmer4):

Actual value of \(\cos 0 = \)

OpenStudy (anonymous):

oh 1-1.26 ? so the difference is 0.3?

OpenStudy (whpalmer4):

I give up. \[\cos 0 = 1\]\[g(0) = 1.26\]\[\cos 0 - g(0) = 1-1.26 = -0.26\]

OpenStudy (anonymous):

oh oops i meant difference is 0.26 so about 0.3 ?

OpenStudy (anonymous):

yes what you said haha :P

OpenStudy (whpalmer4):

I really don't know how I could have been any more clear. What does the expression "take x and subtract y" mean to you?

OpenStudy (whpalmer4):

it means \(x-y\) No, the difference is NOT 0.26!!!

OpenStudy (anonymous):

oh i forgot the - sign.... whoops!!! sorry!! so -0.26, so about -0.3 ? yikes... so sorry!!

OpenStudy (whpalmer4):

Yes, thank you.

OpenStudy (anonymous):

ah sorry again! but thanks so much for your help!!

OpenStudy (whpalmer4):

One of the most valuable lessons to be gained from taking math classes is not the Pythagorean theorem, or the fundamental theorem of calculus, or anything like that. It's the importance of working and reading carefully and methodically.

OpenStudy (anonymous):

yes i agree! i will definitely be reading and working a lot more carefully from now on!! sorry about that! i probably made you really exasperated at my thinking just now haha.. i apologize for that!! but thank you so much for continuing to explain it to me!!

OpenStudy (whpalmer4):

You say that you forgot the minus sign, but I'll bet you help on another problem that you were subtracting 1.26 - 1 because the approximation is higher on the graph than the function...

OpenStudy (whpalmer4):

I don't give up easily when I think the student is capable of understanding what I have to say :-)

OpenStudy (whpalmer4):

But while many would have just told you the answer, I think you really learn it better if you figure it out yourself, even if it raises my blood pressure in the meantime :-)

OpenStudy (whpalmer4):

It's too easy to look at what someone else did and convince yourself that you understand it, but then when they aren't there and you have to do it yourself, suddenly it doesn't seem nearly so clear...

OpenStudy (anonymous):

hahaa i like the bet :P but at first, i did do 1.26-1, but then i did it the x-y thingy... but i thought it couldn't be negative for some odd reason (<--my brain malfunctioned a bit there just now lol) but maybe let's call that bet off? :P hahaa since technically i lost it as i did it incorrectly lol :P thank you:) i really appreciate it!!!sorry about the blood pressure thing haha but i'm really glad you kept explaining it to me :) i think that helps a lot more in the long run :)

OpenStudy (whpalmer4):

If you go look at the 2nd graph I provided again, you'll notice that the error is always negative, except at the point where it just kisses the x-axis at x = 0.707 (pi/4) and there it is 0.

OpenStudy (whpalmer4):

it seems a bit counter-intuitive (and indeed, I had to go verify my recollection) that the error is negative when the approximation came out to be larger than the function, but that's the way it is. I don't doubt that this caused some of your desire to give me positive numbers for the maximum error :-)

OpenStudy (whpalmer4):

and if the problem had said "the magnitude of the error" instead of "the maximum error" your answer would have been fine.